On the basis of following data:
C = 20 + 0.5Y
I = 20
∆ I = 20
(a) Calculate initial level of national income and the change in national income.
(b) Show the multiplier process upto four time periods.
1
Expert's answer
2021-07-16T09:35:30-0400

initial level of income

Y=C+I+G

Y=20+0.5Y+20

Y-0.5Y=20+20

0.5Y/0.5=40/0.5

Y=80

change in national income

Y=20+0.5(80)+20

=80

=20/80x100%

=25%

the multiplier process

1/1-MPC=

change in C/change in Y=20+0.5(80)

=60

change in Y=80

60/80=0.75

MPC=1/1-0.75=1/0.25

=4